# references.bib

### v1

@@ 2.6.3 Continuous stirred tank reactor
with $\mathbf{V}=\textrm{diag}(0.025,1,1)$. The state equation is defined as $f(\mathbf{x},\mathbf{u})=[f_{1}(\mathbf{x},\mathbf{u}),f_{2}(\mathbf{x},\mathbf{u}),f_{3}(\mathbf{x},\mathbf{u})]^{\top}$ with
@@
The policy has to minimise the distance between the temperature in the reactor and the setpoint temperature $T_{r}^{\ast}$, while also minimising the control, i.e. the coolant flowrate. Only a single experiment is performed regarding the CSTR. This experiment tests policies for their generalisability to an industrial environment. The parameters of the CSTR are different in each trial, to simulate as if the policy is operating in multiple scenarios. In total, eight parameters vary between trials (Table 3 in Appendix C). The function set of the GP algorithm is extended to include the exponential and logarithmic functions to aid expressions to deal with the values of the temperatures, which typically range between 200 and 600. The best dynamic symbolic policy evolved with GP is tested in a trial with changing setpoint temperatures, to analyse how the policy and latent state variables respond to these changes. To generate the observations, $\mathbf{D}$ is defined as the identity matrix and $\mathbf{\Sigma}=\textrm{diag}(0.015,3,3)$, where the observation of the concentration is limited to fall within 0 and 1.
@@ 3.5 Generalising symbolic policies are discovered for industrial application
The average best fitness of the NDE policies converges to a worse value than the symbolic policies, which is surprising given the previous experiments (Figure 3G, green curve). A potential reason for the poor performance of the NDE policies could be that they cannot easily approximate the non-linearity required to stabilise the reactor efficiently, which GP can do more easily. The validation fitness of the NDE policies is lower than the symbolic policies, and contain a few bad-performing scores (Figure 3G, green dots).
GP-S converges to a lower fitness during evolution than NDE (Figure 3G, blue curve), and the validation fitness shows high consistency among the policies (Figure 3G, blue dots). GP-D obtains a similar average best fitness as GP-S (Figure 3G, orange curve). The validation fitness of GP-D mostly overlaps with the static policies, but some dynamic symbolic policies obtain worse fitness than any static policy (Figure 3G, orange dots). The best static and dynamic policies are shown in Table 5 (CSTR, both Static and Dynamic). The evolved symbolic policies include non-linear terms such as exponentials and logarithms, but also multiplications of different variables. Such complex non-linear features are more difficult for the NDE to approximate, given the number of neurons in the latent state of the NDE. Increasing the number of neurons would improve its capability to approximate any function, however the optimisation of the network would become less efficient, requiring more computation power. GP can evolve these non-linear functionalities easily, which is yet another advantage of our method.
@@ 3.6 The latent states in a dynamic symbolic policy are transparent
The previous experiments showed that the dynamic symbolic policies are transparent, but the problems of interest are simple control problems. For application purposes, it is more relevant if the policies remain interpretable in an industrial setting, such as the CSTR. The best policy discovered in the experiment conducted on the CSTR is tested in a single trial to analyse the control force and memory capacity (Table 5 CSTR, Dynamic). We interpret the readout of this policy as the sum of both latent variables and a constant. The state equation of variable $a_{2}$ is interpreted to contain an error estimate between the temperature and the setpoint temperature. Furthermore, a self-connection is added to the error estimate to make it less affected by the noisy observations and the error estimate is scaled by the control. The dynamics of variable $a_{1}$ are governed by an exponential of the difference between $a_{2}$ and $a_{1}$.
To increase the number of transparent responses in the trial, the setpoint temperature in the CSTR is changed several times during the trajectory. Therefore the policy has to adjust its control signal to change the temperature accordingly. The policy efficiently stabilises the temperature around each setpoint temperature (Figure 4D, top row). The control signal of the policy heavily changes when a new setpoint temperature is required. The concentration of the reactant in the main reactor goes to zero quickly and only slightly increases when the control input changes (Figure 4D, middle row). Variable $a_{1}$ increases as long as $a_{2}$ is higher, but it converges when $a_{1}$ is substantially higher than $a_{2}$, because the exponential approximates zero (Figure 4b, bottom row, purple curve). Variable $a_{2}$ reacts quickly to the changing setpoint temperatures, but maintains its value once the reactor is stabilised again (Figure 4b, bottom row, blue curve). Variable $a_{1}$ contributes to the control as a constant baseline, which allows variable $a_{2}$ to compute an error-estimate that reacts quickly to changes.
@@ Appendix C Continuous stirred tank reactor parameters
$c$ gmol/liter Reactant concentration $c(0)\sim U(0.4,0.6)$

### v2

@@ 2.6.3 Continuous stirred tank reactor
with $\mathbf{V}=\textrm{diag}(0.025,3,3)$. The state equation is defined as $f(\mathbf{x},\mathbf{u})=[f_{1}(\mathbf{x},\mathbf{u}),f_{2}(\mathbf{x},\mathbf{u}),f_{3}(\mathbf{x},\mathbf{u})]^{\top}$ with
@@
The policy has to minimise the distance between the temperature in the reactor and the setpoint temperature $T_{r}^{\ast}$, while also minimising the control, i.e. the coolant flowrate. Only a single experiment is performed regarding the CSTR. This experiment tests policies for their generalisability to an industrial environment. The parameters of the CSTR are different in each trial, to simulate as if the policy is operating in multiple scenarios. In total, eight parameters vary between trials (Table 3 in Appendix C). Additionally, the concentration in the reactor is not observed, and the observations of the temperature in the reactor and cooling jacket are noisy. The function set of the GP algorithm is extended to include the exponential and logarithmic functions to aid expressions to deal with the values of the temperatures, which typically range between 200 and 600. The best dynamic symbolic policy evolved with GP is tested in a trial with changing setpoint temperatures, to analyse how the policy and latent state variables respond to these changes. To generate the observations, $\mathbf{D}$ is defined such that the concentration is excluded from the observations and $\mathbf{\Sigma}=\textrm{diag}(7.5,7.5)$.
@@ 3.5 Generalising symbolic policies are discovered for industrial application
The average best fitness of the NDE policies has not converged yet after the dedicated number of generations, while being stuck at a poor fitness for the first half of the evolution (Figure 3G, green curve). A potential reason for the lower convergence efficiency of the NDE policies could be that they cannot easily approximate the non-linearity required to stabilise the reactor efficiently, which GP can do more easily, indicated by the faster convergence at a better fitness. The validation fitness of the NDE policies is lower on average than the symbolic policies, and contain a few bad-performing scores (Figure 3G, green dots).
GP-S converges to a lower fitness during evolution than NDE (Figure 3G, blue curve), and the validation fitness shows high consistency among the policies (Figure 3G, blue dots). GP-D obtains a better average best fitness as GP-S (Figure 3G, orange curve). GP-D discovered several policies that outperform GP-S and NDE in terms of validation fitness (Figure 3G, orange dots). However, the high-performing dynamic symbolic policies are again not consistently found, as some policies obtain a validation fitness significantly worse than any static policy.
The best static and dynamic policies are shown in Table 5 (CSTR, both Static and Dynamic). The evolved symbolic policies include non-linear terms such as exponentials and logarithms, but also multiplications of different variables. Such complex non-linear features are more difficult for the NDE to approximate, given the number of neurons in the latent state of the NDE. Increasing the number of neurons would improve its capability to approximate any function, however the optimisation of the network would become less efficient, requiring more computation power. GP can evolve these non-linear functionalities easily, which is yet another advantage of our method.
@@ 3.6 The latent states in a dynamic symbolic policy are transparent
The previous experiments showed that the dynamic symbolic policies are transparent, but the problems of interest are simple control problems. For application purposes, it is more relevant if the policies remain interpretable in an industrial setting, such as the CSTR. The best policy discovered in the experiment conducted on the CSTR is tested in a single trial to analyse the control and memory of the policy (Table 5 CSTR, Dynamic). We interpret the state equation of variable $a_{1}$ as a scaling factor that converges to 2.92 during the trajectory. The state equation of variable $a_{2}$ computes the error between the temperature and setpoint, and adds a self-connection to make the error estimate less affected by noise observations. The error estimate is scaled by the sum of the control, $a_{2}$ itself and a constant. The readout scales $a_{2}$ with the squared of $a_{1}$, and adds a multiplication of $a_{1}$ with the logarithm of the setpoint temperature. The control therefore reacts to differences between the setpoint and current temperature, which is scaled by a factor that increases during the trajectory.
To increase the number of transparent responses in the trial, the setpoint temperature in the CSTR is changed several times during the trajectory. Therefore the policy has to adjust its control signal to change the temperature accordingly. The policy efficiently stabilises the temperature around each setpoint temperature (Figure 4D, top row). The control signal of the policy heavily changes when a new setpoint temperature is required, even remaining close to zero when the temperature has to increase. The concentration of the reactant in the main reactor goes to zero quickly and only slightly increases when the control input changes (Figure 4D, middle row). Variable $a_{1}$ converges to 2.92 as expected (Figure 4D, bottom row, blue curve), while $a_{2}$ shows a large response when the setpoint temperature changes. Together, the latent variables allow the policy to change its control output accordingly.
@@ Appendix C Continuous stirred tank reactor parameters
$c$ gmol/liter Reactant concentration $c(0)\sim U(0.5,1.0)$